OpenStudy (anonymous):

Suppose that y is directly proportional to x. (a) Use the given formula to find the constant proportionality k. (b) Then use y=kx to find y for x=6. y=-68 when x=5

OpenStudy (anonymous):

y is directly proportional to x means that: y=kx for some real number k, to find the value of k, just use the info that y=-68 when x=5.

OpenStudy (anonymous):

(a)k=

OpenStudy (anonymous):

y=kx => k = y/x = -68/5, for x=6 => y=-68 *6 /5 = 81.6

OpenStudy (anonymous):

try to do it, it's so easy.. just substitute y=-68, and x=6 in the equation y=kx.. and solve for k.

OpenStudy (anonymous):

I just need the answer I have a crapload to do. Yes, a crapload! :)

OpenStudy (anonymous):

good luck!

OpenStudy (anonymous):

Thanks so much......."AnwarA".......sigh.

OpenStudy (anonymous):

Nexien, I don't understand what u typed.

OpenStudy (anonymous):

no problem.. but why is the "sigh"? :P

OpenStudy (anonymous):

Why the "good luck!"

OpenStudy (anonymous):

HAHA.. I was just wishing you luck, is there a problem in that?

OpenStudy (anonymous):

Yeah sure.

OpenStudy (anonymous):

Well I am sorry then.